# EMOVA[Uncaptioned image]: Empowering Language Models to See, Hear and Speak with Vivid Emotions

### v2

@@ 4 Omni-modal Alignment and Instruction Tuning
To achieve the omni-model alignment, it is ideal to utilize large-scale omni-modal image-text-speech data, which, however, is either without reach due to copyright (Nagrani et al., 2022) or limited in the quality (Miech et al., 2019). An alternative is to use existing image-text data with TTS-synthesized speeches, which is not only computationally expensive but also hampers data diversity, as most TTS tools generate speech in similar patterns. Recent works (Chen et al., 2024b; Fu et al., 2024b) choose to integrate the speech modality into a well-structured VLLM via a sequential training manner with bi-modal alignment datasets. However, the relationships between different modalities and how to effectively leverage multiple bi-modal alignment datasets remain an open question.
@@
In this work, we explore omni-modal text-centric alignment by utilizing publicly available bi-modal alignment datasets, including both image-text (e.g., captioning) and speech-text (e.g., ASR and TTS) datasets. With the text modality as a bridge, EMOVA ultimately becomes a unified system capable of understanding and generating multiple modalities in a coherent and integrated manner. Specifically, in Sec. 4.1, we explore the following three questions:
@@
  3. 3.
    Is sequential alignment of multiple modalities optimal?
Then we introduce the omni-modal instruction tuning pipeline and the overall training paradigm of our EMOVA in Sec. 4.2 and Sec. 4.3, respectively.
[…]
@@ 4.1 Omni-modal Text-Centric Alignment
For speech capabilities, we directly evaluate the aligned model’s performance on the ASR task of LibriSpeech (Panayotov et al., 2015), while for vision-language, we fine-tune the model using a small amount of high-quality visual instruction data (i.e., the 665K SFT data from ShareGPT4V (Chen et al., 2023d)) and evaluate the fine-tuned model on common vision-language benchmarks. Check Appendix C for evaluation details. Fig. 2 shows the comparison among different paradigms on vision-language (left and middle) and ASR (right, where we report the $1-\text{WER}$ value for better readability) benchmarks, from which we can derive the following observations:
@@
We find 1) Joint outperforms Joint-entangled on vision-language benchmarks, and 2) in the speech tasks, Joint maintains significant advantages over its entangled counterpart. This can be attributed to the semantic-acoustic disentanglement which makes speech units more analogous to languages, a domain LLMs are particularly specialized.
@@
In light of these observations, we have chosen to pursue the ultimate alignment strategy that simultaneously aligns image-text and speech-unit-text for EMOVA, which offers two important benefits, 1) it fosters mutual enhancement between vision-language and speech capabilities, and 2) it avoids the issue of catastrophic forgetting during sequential alignment of multiple modalities.
Figure 3: Demonstration of the omni-modal instruction tuning. 1) To empower emotional spoken dialogues, EMOVA is trained to explicitly select the speech style labels (e.g., emotions and pitches) with output speech units. 2) For the ease of parsing, data components are arranged in JSON format.
@@ 4.2 Omni-modal Instruction Tuning
After the omni-modal text-centric alignment, the model learns fundamental vision-language (e.g., captioning) and speech capabilities (e.g., ASR and TTS). However, instruction tuning is essential to better follow complicated user instructions and respond with vivid emotions.
@@
Data organization and chain of modality.
The omni-modal instruction data can be represented as $D_{\text{omni}}=\{(x_{V},u_{S},x^{o}_{T},c^{o}_{\text{style}},u^{o}_{S})_{i}\}_{i=1}^{N}$, where the input consists of the optional queried image $x_{V}$ and the speech units of the instruction $u_{S}$, while the output consists of the textual response $x^{o}_{T}$, the predicted speech style labels $c^{o}_{\text{style}}$, and the output speech unit $u^{o}_{S}$. Note that we train EMOVA to explicitly select styles (e.g., emotions and pitches), which are utilized to determine the corresponding style embedding for the U2S detokenizer (Sec. 3.2). Moreover, since directly generating the speech responses is challenging, we decompose the speech response procedure into three primary steps: 1) recognizing user instructions into texts; 2) generating textual responses based on the recognized instructions; 3) generating the style labels and response speech units based on the textual responses. For ease of parsing during deployment, the target outputs are formatted as JSON, as shown in Fig. 3.
Settings Stage-1 Stage-2 Stage-3
@@
# Tokens Max $256\times(1+9)$ Max $256\times(1+9)$ Max $256\times(1+9)$
Data Dataset LCS EMOVA-Alignment (Fig. 10) EMOVA-SFT (Fig. 4)
# Samples 558K 7.4M 4.4M
Training Trainable Projector Full Model (Half ViT) Full Model
Batch Size 256 256 128
LR: $p(\cdot)$ $1\times 10^{-3}$ $2\times 10^{-5}$ $2\times 10^{-5}$
LR: $v(\cdot)$ $2\times 10^{-5}$ $2\times 10^{-6}$
LR: $f(\cdot)$ $2\times 10^{-5}$ $2\times 10^{-5}$
Epoch 1 1 1
@@ 4.3 Overall Training Paradigm
Inspired by Chen et al. (2023d), a three-stage training paradigm is adopted for EMOVA,
  • Stage-1: Vision-language pre-alignment. The purpose is to align visual features into the embedding space of LLMs. Only the vision projector $p(\cdot)$ is trained.
  • Stage-2: Omni-modal text-centric alignment. This stage jointly performs the vision-language and speech-language alignment simultaneously. We train the LLM $f(\cdot)$, vision projector $p(\cdot)$, and the deeper half of vision encoder $v(\cdot)$ layers.
  • Stage-3: Omni-modal instruction tuning. To empower EMOVA to respond accurately to omni-modal instructions, we organize different datasets with various types of instructions to enforce EMOVA to learn generalization across tasks, as detailed in Sec. 5.1.
Table 3: Comparison on vision-language and speech benchmarks. 1) EMOVA surpasses GPT-4V and Gemini Pro 1.5 on 10 of 14 vision-language benchmarks, while reaching over 95% of GPT-4o performance on nearly all benchmarks. 2) Meanwhile, EMOVA achieves state-of-the-art performance on the ASR task, surpassing its speech counterparts by a significant margin.
Benchmarks EMOVA Gemini GPT- GPT- LLaVA- Intern- Mini- AnyGPT VITA
8B Pro 1.5 4V 4o OV-7B VL2-8B Omni 7B 8x7B
MME 2205 - 1927 2310 1998 2215 - - 2097
MMBench 82.8 - 75.0 83.4 80.8 81.7 - - 71.8
SEED-Image 78.1 - 71.6 77.1 - 75.4 - - -
MM-Vet 55.8 - 67.7 - 57.5 54.3 - - 41.6
RealWorldQA 64.3 68.7 61.4 75.4 66.3 - - - -
TextVQA 82.0 73.5 77.4 - - 77.4 - - -
ChartQA 81.8 81.3 78.5 85.7 80.0 83.3 - - -
DocVQA 90.4 86.5 88.4 92.8 87.5 91.6 - - -
InfoVQA 64.4 72.7 - - 68.8 74.8 - - -
OCRBench 824 - 656 736 - 794 - - 678
MathVista 61.1 52.1 49.9 63.8 63.2 58.3 - - 44.9
Mathverse 27.8 - 33.6 - 26.2 - - - -
ScienceQA-Img 94.0 - 75.7 - 96.0 97.1 - - -
AI2D 82.8 80.3 78.2 84.6 - 83.8 - - 73.1
[…]
@@ 5.1 Training configuration
In this stage, we only train the parameters of the vision projector $p(\cdot)$ for vision-language pre-alignment with the LCS-558K dataset (Liu et al., 2024b), with the high-resolution image-slicing strategy (Liu et al., 2024a) adopted.
@@
We assemble a unified dataset with 7.4M samples for both the image-text and speech-text alignment, as summarized in Fig. 10. Specifically, we utilize pre-training datasets from ShareGPT4V (Chen et al., 2023d), ALLaVA (Chen et al., 2024a) (both the original English version and the Chinese version translated on our own), and ShareGPT-4o (Cui et al., 2023) for general perception, while for the OCR capabilities, we leverage SynthDog (Kim et al., 2022), MMC-Alignment (Liu et al., 2023a), K12 Printing, and UReader Text Reading subset (Ye et al., 2023). Moreover, we use the 2,000 hours of ASR and TTS data from LibriSpeech (Panayotov et al., 2015) and AISHELL-2 (Du et al., 2018) for speech-text alignment, and to preserve the language capabilities of LLMs, we further incorporate the text-only data from Magpie Pro (Xu et al., 2024).
@@
We collect the EMOVA-SFT dataset consisting of 4.4M multi-task omni-modal samples (see Fig. 4). We start by gathering high-quality open-sourced visual instruction datasets, including ShareGPT4V (Chen et al., 2023d), InternVL (Chen et al., 2024b), Meteor (Lee et al., 2024), Idefics-2 (Laurençon et al., 2024), Cambrian (Tong et al., 2024), and LLaVA-Onevision (Li et al., 2024a), followed by quality checking, re-formatting all data samples to a unified template, and removing the duplicates. For speech, we include the training split of EMOVA omni-model instruction data (c.f., Sec. 4.2), with 10% of speech alignment datasets to maintain ASR and TTS performance. We adopt 128 Ascend 910 Training NPUs in parallel for model training (check more details in Table 2).
5.2 Comparison to the SOTA Models
Experimental results are provided in Table 3. We compare a wide range of state-of-the-art VLLMs, including proprietary ones like Gemini Pro 1.5 (Reid et al., 2024), GPT-4V (OpenAI, 2023), GPT-4o (OpenAI, 2024), and open-sourced ones like LLaVA-OneVision-7B (Li et al., 2024a), InternVL2-8B (Chen et al., 2024b), together with the Speech LLM Mini-Omni (Xie & Wu, 2024) and omni-modal LLM AnyGPT (Zhan et al., 2024) and VITA-8x7B (Fu et al., 2024b).
@@ 5.2 Comparison to the SOTA Models
As an omni-modal model, EMOVA achieves comparable performance with state-of-the-art open-sourced VLLMs across multiple benchmarks. Specifically, our EMOVA outperforms both InternVL2 and LLaVA-OV on MMBench, SEED-Image, TextVQA, and OCRBench consistently, while demonstrating exceptional proficiency in solving math problems requiring precise visual content interpretation. EMOVA achieves a 2.8% improvement on MathVista compared with InternVL2 and 1.6% higher score on Mathverse compared to LLaVA-OneVision.
Furthermore, EMOVA shows competitive performance even compared with the proprietary models. EMOVA exceeds both GPT-4V and Gemini Pro 1.5 significantly on 10 out of 14 benchmarks, while for GPT-4o, EMOVA outperforms on both SEEDBench-Image and OCRBench, reaching over 95% of GPT-4o’s performance on ALL evaluated benchmarks except RealWorldQA.
@@
Compared with VITA-8x7B, EMOVA shows substantial improvement on visual-language benchmarks. Specifically, EMOVA is 112 points higher than VITA on MME, and surpasses VITA by 21.5% on OCRBench (824 vs. 678), underscoring the effectiveness of our approach and the potential to push boundaries of omni-modal abilities. What’s more, ASR results show that EMOVA significantly outperforms the most recent omni-modal model VITA, even surpassing its SLLM counterpart Mini-Omni, showing the effectiveness of the semantic-acoustic disentanglement and omni-modal mutual benefits. We further report text-to-speech (TTS) results in Table 6. See Fig. 5, Fig. 6 and Fig. 8 for qualitative examples. More results can be found in Appendix G. For the first time, our EMOVA obtains state-of-the-art performance on both the vision-language and speech benchmarks simultaneously.
@@ 5.3 Evaluation of Emotion-Rich Spoken Dialogue
In this section, we evaluate the end-to-end spoken dialogue capabilities of EMOVA. As discussed in Sec. 4.2, the model takes an input image $x_{V}$ and user instructions in the form of speech units $u_{S}$, and outputs text responses, style labels, and corresponding speech units. To ensure comprehensive evaluation, we propose the following evaluation metrics:
@@
    Unit-input-text-output score focuses on the quality of the text responses of LLM when the inputs are speech units, bypassing errors from speech synthesis.
@@
    Text-input-text-output score inputs the ground-truth text of the user instruction and evaluate the model’s text output. This helps disentangle the impact of speech recognition errors and eliminates the influence of the JSON format.
@@
    ASR and TTS evaluate how accurately EMOVA recognizes speech units and how effectively it generates speech units from text. See Appendix D for more details.
@@
    Style label classification accuracy evaluates the accuracy of the model in selecting the appropriate speech style labels (Sec. 3.2).
@@
    Style controllablity assesses the controllability of U2S detokenizer with the given conditional style labels using the confusion matrix that compares the generated and recognized style labels. See Appendix D for more details.
Due to the lack of emotionally rich spoken dialogue evaluation datasets, we split a test set from our synthesized omni-modal instruction-tuning data (see Sec. 4.1). GPT-4o are used for automated evaluation. Details are provided in Appendix D.
@@ Results.
Table 4 the spoken dialogue performance of EMOVA.
(i) By comparing the end-to-end dialogue score with the unit-input-text-output score, we notice that the two scores are closely aligned, with a maximum gap of only $0.22$, except for Speech-Image-ZH. The TTS-WER/CER is generally low, revealing that EMOVA can synthesize accurate speech based on text responses. However, the Speech-Image-ZH is an outlier, which we attribute to its complexity. It includes tasks such as generating poetries and answering riddles, resulting in more intricate responses. When these answers are converted to speeches and then transcribed back to texts, multiple variations often arise, leading to discrepancies from the original responses.
(ii) Comparing the unit-input-text-output score with the text-input-text-output score, we observe that their differences correlate with the ASR results of speech instructions. Specifically, for Speech-Text-EN and Speech-Text-ZH, which involve more complex instructions,
Figure 7: Confusion matrix between the generated and recognized emotions. The emotions generated by our U2S detokenizer are recognized with high probability.
EMOVA reports inferior ASR performance ($6.8$ and $12.6$, respectively) compared to other datasets ($2.45$ and $1.85$). Consequently, when we replace speech instructions with ground-truth transcriptions, EMOVA shows significant improvements from unit-input to text input score. On the contrary, for datasets with accurate ASR performance, the results are quite similar, suggesting EMOVA retains robust dialogue capabilities when using the JSON format.
(iii) Examining the classification accuracy of style labels, we find that EMOVA performs satisfactorily in classifying emotions and pitch during speech conversations, achieving an accuracy of over 80%. The confusion matrix comparing the conditional and recognized emotion labels is shown in Fig. 7. The results indicate that the four emotions are recognized with high probabilities, with three achieving over 80% accuracy. This demonstrates that our U2S detokenizer effectively controls common emotions, endowing the synthesized speech with vivid emotional expression.
Table 4: The performance of EMOVA on Speech Dialogue tasks. By default, we evaluate on the corresponding test set of the evaluated datasets.
[…]
Speech-Image-EN 7.21 7.43 7.75 89.01 98.30 2.45 4.60
Speech-Text-EN 6.90 7.02 8.27 87.26 90.50 6.80 6.75
Speech-Image-ZH 5.87 6.51 6.44 92.60 96.60 1.85 13.95
Speech-Text-ZH 4.70 4.88 6.00 87.19 94.40 12.60 9.70
Figure 8: EMOVA engages in spoken dialogue with a cheerful tone.

### v3

@@ 4 Training Omni-modal LLMs
To achieve omni-model alignment, it is ideal to use large-scale omni-modal image-text-speech data, which, however, is either without reach due to copyrights [67] or limited in quality [65]. An alternative is to use the existing image-text data with the TTS-synthesized speeches, which is not only computationally expensive but also hampers data diversity, as most TTS tools generate speeches in similar patterns. Recent works [10, 19] choose to integrate the speech modality into a well-structured VLLM via a sequential training manner with bi-modal alignment data. However, the relationships among modalities and how to effectively leverage multiple bi-modal alignment datasets remain unclear.
@@
In this work, we explore omni-modal text-centric alignment by utilizing the publicly available bi-modal alignment datasets, including image-text (e.g., image captioning) and speech-text (e.g., ASR and TTS) datasets. With text modality as a bridge, our EMOVA ultimately becomes a unified system capable of understanding and generating multiple modalities in a coherent and integrated manner. In Sec. 4.1, we first explore the following three questions:
@@
We then discuss the omni-modal instruction tuning pipeline and the overall training paradigm of EMOVA in Sec. 4.2 and Sec. 4.3, respectively.
@@ 4.1 Omni-modal Text-Centric Alignment
For speech abilities, we evaluate the aligned model’s performance on the ASR task of LibriSpeech [72], while for the vision-language, we fine-tune the model with a small amount of high-quality visual instruction data (i.e., 665K SFT data from the ShareGPT4V [9]) and evaluate the fine-tuned model on common vision-language benchmarks. Check Appendix C for evaluation details. Fig. 3 shows the comparison among different paradigms on vision-language (left and middle) and ASR (right, where we report the $1-\text{WER}$ value for better readability) benchmarks, from which we can derive the following observations:
@@
We find that 1) Joint outperforms Joint-entangled on vision-language benchmarks, and 2) in the speech tasks, Joint maintains significant advantages over its entangled counterpart. This can be attributed to the semantic-acoustic disentanglement which makes speech units more analogous to languages.
Benchmarks EMOVA EMOVA EMOVA Gemini GPT- GPT- Whisper Mini- VITA VITA Baichuan-
3B 7B 72B Pro 1.5 4V 4o Large Omni2 8x7B 1.5 Omni-7B
MME 2175 2317 2402 - 1927 2310 - - 2097 2311 2187
MMBench 79.2 83.0 86.4 - 75.0 83.4 - - 71.8 76.6 76.2
SEED-Image 74.9 75.5 76.6 - 71.6 77.1 - - 72.6 74.2 74.1
MM-Vet 57.3 59.4 64.8 - 67.7 - - - 41.6 51.1 65.4
RealWorldQA 62.6 67.5 71.0 68.7 61.4 75.4 - - 59.0 66.8 62.6
TextVQA 77.2 78.0 81.4 73.5 77.4 - - - 71.8 74.9 74.3
ChartQA 81.5 84.9 88.7 81.3 78.5 85.7 - - 76.6 79.6 79.6
DocVQA (test) 93.5 94.2 95.9 86.5 88.4 92.8 - - - - -
InfoVQA (test) 71.2 75.1 83.2 72.7 - - - - - - -
OCRBench 803 814 843 - 656 736 - - 678 752 700
ScienceQA-Img 92.7 96.4 98.2 - 75.7 - - - - - -
AI2D 78.6 81.7 85.8 80.3 78.2 84.6 - - 73.1 79.3 -
MathVista 62.6 65.5 69.9 52.1 49.9 63.8 - - 44.9 66.2 51.9
Mathverse 31.4 40.9 50.0 - 33.6 - - - - - -
[…]
Table 2: Comparison on vision-language and speech benchmarks. 1) EMOVA outperforms GPT-4o/4V and Gemini Pro 1.5 on 11 of the 14 vision-language benchmarks, providing a powerful open-sourced alternative. 2) Meanwhile, our EMOVA achieves state-of-the-art performance on Librispeech, surpassing its speech and omni-modal counterparts significantly. : reported by [43].
@@ 4.2 Omni-modal Instruction Tuning
After the omni-modal text-centric alignment in Sec. 4.1, the model learns the fundamental vision-language (e.g., captioning) and speech capabilities (e.g., ASR and TTS). However, instruction tuning is essential to better follow complicated user instructions and respond with vivid emotions.
@@
Data organization and the chain of modality.
The omni-modal instruction data can be represented as $D_{\text{omni}}=\{(x_{V},u_{S},x^{o}_{T},c^{o}_{\text{style}},u^{o}_{S})_{i}\}_{i=1}^{N}$, where the input consists of the optional queried image $x_{V}$ and the speech units of the instruction $u_{S}$, while the output consists of the textual response $x^{o}_{T}$, the predicted speech style labels $c^{o}_{\text{style}}$, and the output speech unit $u^{o}_{S}$. Note that we train EMOVA to explicitly select styles (e.g., emotions and pitches), which are utilized to determine the corresponding style embedding for the U2S detokenizer (Sec. 3.2). Furthermore, since directly generating speech responses is challenging, we decompose the speech response procedure into three primary steps: 1) recognizing user instructions into texts; 2) generating textual responses based on the recognized instructions; 3) generating the style labels and response speech units based on the textual responses. For ease of parsing during deployment, the target outputs are formatted as JSON, as in Fig. 4.
@@
Speech-Image-EN 7.45 7.56 7.95 82.50 97.70 2.40 3.20
Speech-Text-EN 6.85 6.90 7.38 81.20 84.70 6.90 2.90
Speech-Image-ZH 6.48 7.02 6.82 77.60 95.90 1.70 12.00
Speech-Text-ZH 5.25 5.58 6.60 80.90 93.20 10.70 12.20
Table 3: Evaluation of EMOVA-7B on Speech Dialogue. By default, we evaluate on the corresponding test set of the evaluated datasets.
@@ 4.3 Overall Training Paradigm
Inspired by [9], a three-stage training paradigm is adopted,
  • Stage-1: Vision-language pre-alignment. The purpose is to align the visual features into the embedding space of LLMs. Only the vision projector $p(\cdot)$ is trained.
  • Stage-2: Omni-modal text-centric alignment. This stage performs vision-language and speech-language alignment jointly. We train the LLM $f(\cdot)$, vision projector $p(\cdot)$, and the deeper half of vision encoder $v(\cdot)$ layers.
  • Stage-3: Omni-modal instruction tuning. We organize different datasets with various types of instructions to learn generalization across tasks, as detailed in Sec. 5.1.
@@ 5.1 Training configuration
In this stage, we only train the parameters of the vision projector $p(\cdot)$ for vision-language pre-alignment with the LCS-558K dataset [49], with the high-resolution image-slicing strategy [48] adopted.
@@
We assemble a unified dataset with 7.4M samples for both the image-text and speech-text alignment, as summarized in Fig. 9. Specifically, we utilize pre-training datasets from ShareGPT4V [9], ALLaVA [4] (both the original English version and the Chinese version translated on our own), and ShareGPT-4o [12] for general perception, while for the OCR capabilities, we leverage SynthDog [34], MMC-Alignment [47], K12 Printing, and the UReader Text Reading subset [89]. Moreover, we adopt the 2,000 hours of ASR and TTS data from LibriSpeech [72] and AISHELL-2 [15] for speech-text alignment, and to preserve the language capabilities of LLMs, we further incorporate the text-only data from Magpie Pro [88]. Check more details in Fig. 9.
@@
We collect the EMOVA-SFT dataset consisting of 4.4M multi-task omni-modal samples (see Fig. 8). We start by gathering high-quality open-sourced visual instruction datasets, including ShareGPT4V [9], InternVL [10], Meteor [38], Idefics-2 [36], Cambrian [80], and LLaVA-Onevision [40], followed by quality checking, re-formatting all data samples with a unified template, and removing the duplicated data. For speech, we include the training split of EMOVA omni-model instruction data (cf., Sec. 4.2), with 10% of speech alignment datasets to maintain ASR and TTS performance. We train with 128 Ascend 910B (64GB) NPUs in parallel (check more details in Table 5).
5.2 Comparison to SOTA Models
Experimental results are provided in Table 2. We compare a wide range of state-of-the-art VLLMs, including Gemini Pro 1.5 [75], GPT-4V [69], GPT-4o [70], together with the Speech LLM (i.e., Mini-Omni2 [87]) together with the ASR expert Whisper-Large [74], and the omni-modal LLMs (i.e., VITA-8x7B [19], VITA-1.5 [20] and Baichuan-Omni [43]).
@@ 5.2 Comparison to SOTA Models
As an omni-modal model, EMOVA obtains comparable performance with the state-of-the-art VLLMs on multiple vision-language benchmarks, while showing superior proficiency in solving math problems needing precise visual content interpretation. Our EMOVA-7B surpasses GPT-4V by +7.3 on MathVerse, and our EMOVA-72B exceeds GPT-4o by +6.1 on MathVista. On 11 out of the 14 benchmarks, our EMOVA outperforms both GPT-4o/4V and Gemini Pro 1.5, providing a powerful open-sourced alternative.
@@
Under the similar 7B capacity, EMOVA-7B is 220 points higher than VITA on MME, surpassing VITA by 20.6% on OCRBench (814 vs. 678). Moreover, it surpasses Baichuan-Omni-7B, a more recent Omni-modal LLM, on nearly all the evaluated benchmarks, underscoring the effectiveness of our approach and potential to push the boundary of omni-modal abilities.
Moreover, ASR results show that EMOVA significantly outperforms the most recent omni-modal model VITA, even surpassing its SLLM counterparts Mini-Omni2 and ASR-specialized Whisper-Large, showing the effectiveness of the semantic-acoustic disentanglement and omni-modal mutual benefits. We further report TTS results in Table 6, and provide the qualitative examples in Fig. 5 and Appendix G. For the first time, EMOVA obtains state-of-the-art results on both the speech and vision-language benchmarks.
@@ 5.3 Evaluation of Emotion-Rich Spoken Dialogue
In this section, we evaluate the end-to-end spoken dialogue abilities of EMOVA-7B. As stated in Sec. 4.2, the model takes an input image $x_{V}$ and user instructions in the form of speech units $u_{S}$, and outputs text responses, style labels, and corresponding speech units. To ensure comprehensive evaluation, we propose the following evaluation metrics:
@@
    Unit-input-text-output score focuses on the quality of the textual responses of LLM when the inputs are speech units, bypassing errors from speech synthesis.
@@
    Text-input-text-output score inputs the ground-truth user instruction texts and evaluates the model’s text outputs. This helps disentangle the impact of speech recognition errors and eliminates the effect of JSON format.
@@
    ASR and TTS evaluate how accurately the model recognizes the speech units and how effectively it generates speech units from text. See Appendix D for more details.
@@
    Style label classification accuracy evaluates the accuracy in selecting the appropriate speech style labels.
@@
    Style controllablity assesses the controllability of U2S detokenizer with the given conditional style labels using the confusion matrix comparing the generated and recognized style labels. See Appendix D for more details.
Due to the lack of emotionally rich spoken dialogue evaluation datasets, we split a test set from our synthesized omni-modal instruction-tuning data (Sec. 4.1). GPT-4o is used for automated evaluation. See details in Appendix D.
@@ Results.
Table 3 the spoken dialogue performance.
(i) By comparing the end-to-end dialogue score with the unit-input-text-output score, we notice that the two scores are closely aligned, with a maximum gap of only $0.33$, except for Speech-Image-ZH. TTS-WER/CER is generally low for English, revealing that EMOVA can synthesize accurate speech based on textual responses, which, however, is harder for Chinese, which we attribute to its complexity. It includes tasks such as generating poetries and answering riddles, resulting in more intricate responses.
Figure 6: Confusion matrix between the generated and recognized emotions. The emotions generated by our U2S detokenizer are recognized with high probability. Best view with zooming in.
(ii) Comparing the unit-input-text-output score with the text-input-text-output score, we notice that their differences correlate with the ASR performance of the speech instructions, especially for Speech-Text-EN and Speech-Text-ZH, which involve more complex instructions.
Our EMOVA-7B reports inferior ASR performance ($6.9$ and $10.7$, respectively) compared to other datasets ($2.4$ and $1.7$). Consequently, when we replace speech instructions with ground-truth transcriptions, EMOVA shows significant improvements from unit-input to text input score. On the contrary, for datasets with accurate ASR performance, the results are quite similar, suggesting EMOVA retains robust dialogue capabilities when using the JSON format.
(iii) Examining the classification accuracy of style labels, we find that EMOVA performs satisfactorily in classifying emotions and pitches during speech conversations, achieving an accuracy of over 75%. The confusion matrix comparing the conditional and recognized emotion labels is shown in Fig. 6. The results indicate that the four emotions are recognized with high probabilities, with three achieving over 80% accuracy. This demonstrates that our U2S detokenizer effectively controls common emotions, endowing the synthesized speech with vivid emotional expression.